# Approximate contraction of arbitrary tensor networks with a flexible and efficient density matrix algorithm

### v1

@@ 8 Experimental results
In Section 8.3, we justify the partitioned_contract algorithm presented in Algorithm 2. We justify our embedding tree selection algorithm and explore the impact of the environment size on accuracy and efficiency across multiple problems. Additionally, we conduct a comprehensive comparison between the MPS and the comb ansatz. Furthermore, we evaluate partitioned_contract, the CATN algorithm [61] [^1] We use the CATN implementation at https://github.com/panzhang83/catn., and SweepContractor [14] [^2] We use the SweepContractor implementation at https://github.com/chubbc/SweepContractor.jl., in contracting tensor networks defined on lattices and random regular graphs as well as tensor networks from random quantum circuit simulation. We demonstrate a 9.2X speed-up while maintaining the same level of accuracy when contracting tensor networks defined on 3D lattices using the Ising model.
@@ 8.1 Implementations, tested tensor networks, and the evaluation
The proposed algorithms in the paper are being developed at https://github.com/ITensor/ITensorNetworks.jl. ITensorNetworks.jl is a publicly available Julia [6] package built for manipulating tensor networks of arbitrary geometry, and is built on top of ITensors.jl [24]. The library also provides an interface to OMEinsumContractionOrders.jl [^3] The library is implemented at https://github.com/TensorBFS/OMEinsumContractionOrders.jl, which implements multiple heuristics introduced in [27, 44] to generate efficient contraction paths for exact tensor network contractions. For all the results presented in this work, we use the Simulated Annealing bipartition + Greedy algorithm (SABipartite) [44] to generate contraction paths for exact tensor network contractions.
@@ 8.2 Comparison between the density matrix algorithm and the canonicalization-based algorithm
In Fig. 16, we conduct a performance comparison of truncating both MPSs and balanced binary tree tensor networks. Let $R$ denote the rank of the input MPS and the balanced binary tree, the analytical asymptotic cost for truncating an MPS is $\Theta(R^{3})$, whereas for truncating a balanced binary tree is $\Theta(R^{4})$. As depicted in the results, the scaling behavior of both algorithms aligns with the analytical predictions. Despite the density matrix algorithm incurring a constant overhead in terms of GFlops, we observe that it exhibits slightly faster performance. This advantage can be attributed to the fact that the majority of the density matrix algorithm’s execution time is spent on tensor contractions, which are practically faster compared to matrix factorizations, even though both operations have a similar computational complexity.
@@ 8.3 Benchmark of the partitioned_contract algorithm
(a) Ising Model, $\beta=0.3$
(b) Ising Model, $\beta=0.3$
(c) Ising Model, $\beta=0.3$
(d) Random Model, $\alpha=-0.4$
(e) Random Model, $\alpha=-0.4$
(f) Random Model, $\alpha=-0.4$
Figure 20: Performance comparison between partitioned_contract, SweepContractor [14], and CATN [61] in contracting $5\times 5\times 5$ 3D lattices. The swap batch size is fixed to be 32 for all experiments. In the legends, “PC” denotes partitioned_contract, MPS/Comb denotes the embedding tree ansatz, and the values (1, 3, 5) denote the size of each partition. The number shown on top of each point is the maximum bond size $\chi$. In CATN, “Dmax” is an additional input parameter of the algorithm that controls the size of the MPS uncontracted modes.
@@
Figure 21: Performance comparison between partitioned_contract and SweepContractor [14] in contracting random regular graphs with degree 3 and 220 vertices. The swap batch size is fixed to be 32 for all experiments. In the legends, “PC” denotes partitioned_contract, MPS/Comb denotes the embedding tree ansatz, and the number (1, 3, 6) denotes the size of each partition. The number shown on top of each point is the maximum contracted bond size $\chi$.
@@
We explore the impact of the environment size on the accuracy and efficiency of contracting tensor networks defined on 3D lattices and random regular graphs, and the results are shown in Fig. 20 and Fig. 21.
In both 3D lattices and random regular graphs, we employ the maximally-unbalanced partial contraction path for the contraction process. This path initiates from one partition and progressively combines the previously-contracted section with a new partition following a linear sequence of the partitions. For 3D lattices, each partition represents either a portion of a fiber or an entire fiber of the lattice. The contraction path is determined through a row- or column-wise traversal of the 2D array resulting from partitioning the 3D lattice into fibers. Regarding random regular graphs, we draw inspiration from [34] to construct the contraction path using a linear ordering of vertices. We achieve this by first employing recursive bisection to generate the linear ordering of all the vertices. Then, we sequentially include a partition consisting of a specified number of tensors into the contraction path, following the order of traversal in the vertex ordering.
When considering 3D lattices with the MPS ansatz, the results presented in Fig. 20b and Fig. 20e reveal that employing a partition size of 3 or 5 leads to both faster and more accurate contractions when compared to the base condition where each partition contains only one tensor. The improved efficiency arises from using larger partitions, which reduces the number of executions of the density matrix algorithm, offsetting any overhead from using larger environments. Regarding accuracy, we can see that under the same maximum contracted bond size, utilizing a partition size of 3 or 5 yields lower relative errors compared to using a partition size of 1. This observation validates the efficacy of the environment in enhancing accuracy. For the comb ansatz, Fig. 20c and Fig. 20f show that employing a partition size of 3 results in the lowest running time. Similarly to the MPS ansatz, using a partition size of 3 or 5 exhibits better accuracy compared to a partition size of 1. See the next section for a comparison of the MPS and comb structures.
Regarding random regular graphs, the results displayed in Figs. 21b, 21c, 21e and 21f indicate that using a partition size of 6 results in the best combination of efficiency and accuracy. To summarize, employing a larger partition leads to a larger environment size, generally reducing the contraction error under the same rank. However, when it comes to efficiency, the optimal partition size depends on the specific problem. Factors such as the number of executions of the density matrix algorithm to be performed and the cost of forming the density matrix under different environment sizes need to be taken into consideration to determine the most suitable partition size.
@@
In this section we discuss the relative merits of using the MPS and comb tree ansatzes. When contracting 3D lattices, the results in Figs. 20a and 20d demonstrate that both MPS and comb structures exhibit similar performance when the maximum bond size is small. However, as the maximum bond size increases, using the comb ansatz becomes slower in comparison to MPS. On the other hand, when contracting random regular graphs, the results in Figs. 21a and 21d reveal that both structures display similar levels of accuracy and efficiency. In summary, both MPS and comb binary tree structures perform similarly in terms of accuracy. However, the efficiency of the comb structure may lag behind MPS, particularly when dealing with a large rank. This disparity in performance is attributed to the presence of large tensors with size $\chi^{3}$ in the comb ansatz.
@@
Here we compare our proposed partitioned_contract algorithm, along with the CATN algorithm [61] and SweepContractor [14], on contracting tensor networks defined on lattices and random regular graphs. As shown from the results in Figs. 20a and 21a, our algorithm outperforms both CATN and SweepContractor in terms of efficiency across all relative errors. Notably, when contracting a 3D lattice tensor network based on the Ising model, partitioned_contract achieves a 9.2X speed-up compared to both CATN and SweepContractor when reaching a relative error of less than $10^{-9}$. Similarly, when contracting a tensor network with a random regular graph structure based on the Ising model, partitioned_contract achieves a 52.4X speed-up compared to SweepContractor when achieving a relative error of less than $10^{-5}$.

### v3

@@ 1.1 Our contributions
1.2 Organization
This paper is organized as follows. In Sections 2 and 3, we introduce the definitions, the computational cost model, and background for the proposed algorithm. Section 4 provides an overview of the proposed approximate tensor network contraction algorithm, with detailed components discussed in Sections 5, 6 and 7. In Section 8, we present the results of a series of experiments to evaluate the performance of the proposed algorithm.
@@ 8 Experimental results
In Section 8.3, we justify the partitioned_contract algorithm presented in Algorithm 2. We justify our embedding tree selection algorithm and explore the impact of the environment size on accuracy and efficiency across multiple problems. Additionally, we conduct a comprehensive comparison between the MPS and the comb ansatz. Furthermore, we evaluate partitioned_contract, the CATN algorithm [61] [^1] We use the CATN implementation at https://github.com/panzhang83/catn., SweepContractor [14] [^2] We use the SweepContractor implementation at https://github.com/chubbc/SweepContractor.jl., and hyperoptimized approximate contraction [26] [^3] We use the hyperoptimized approximate contraction implementation at quimb [25] (https://github.com/jcmgray/quimb)., in contracting tensor networks defined on lattices and random regular graphs as well as tensor networks from random quantum circuit simulation. We demonstrate a 9.2X speed-up while maintaining the same level of accuracy when contracting tensor networks defined on 3D lattices using the Ising model.
@@ 8.1 Implementations, tested tensor networks, and the evaluation
The proposed algorithms in the paper are being developed at https://github.com/ITensor/ITensorNetworks.jl. ITensorNetworks.jl is a publicly available Julia [6] package built for manipulating tensor networks of arbitrary geometry, and is built on top of ITensors.jl [24]. The library also provides an interface to OMEinsumContractionOrders.jl [^4] The library is implemented at https://github.com/TensorBFS/OMEinsumContractionOrders.jl., which implements multiple heuristics introduced in [27, 44] to generate efficient contraction paths for exact tensor network contractions. For all the results presented in this work, we use the Simulated Annealing bipartition + Greedy algorithm (SABipartite) [44] to generate contraction paths for exact tensor network contractions.
@@ 8.2 Comparison between the density matrix algorithm and the canonicalization-based algorithm
In Fig. 16, we conduct a performance comparison of truncating both MPSs and balanced binary tree tensor networks. Let $R$ denote the rank of the input MPS and the balanced binary tree, the analytical asymptotic cost for truncating an MPS is $\Theta(R^{3})$, whereas for truncating a balanced binary tree is $\Theta(R^{4})$. As depicted in the results, the scaling behavior of both algorithms aligns with the analytical predictions. Despite the density matrix algorithm incurring a constant overhead in terms of GFlops, we observe that it exhibits slightly faster performance. This advantage can be attributed to the fact that the majority of the density matrix algorithm’s execution time is spent on tensor contractions, which are practically faster compared to matrix factorizations, even though both operations have a similar computational complexity. This property makes the density matrix algorithm more favorable on GPUs due to their ability to efficiently run tensor contractions in parallel.
@@ 8.3 Benchmark of the partitioned_contract algorithm
(a) $5\times 5\times 5$, $\beta=0.3$
(b) $6\times 6\times 6$, $\beta=0.3$
(c) $28\times 28$, $\beta=0.44$
(d) $5\times 5\times 5$, $\beta=0.3$
(e) $5\times 5\times 5$, $\beta=0.3$
(f) $5\times 5\times 5$, $\beta=0.3$
(g) $5\times 5\times 5$, $\beta=0.3$
(h) $6\times 6\times 6$, $\beta=0.3$
(i) $6\times 6\times 6$, $\beta=0.3$
(j) $6\times 6\times 6$, $\beta=0.3$
(k) $6\times 6\times 6$, $\beta=0.3$
(l) $28\times 28$, $\beta=0.44$
(m) $28\times 28$, $\beta=0.44$
(n) $28\times 28$, $\beta=0.44$
(o) $28\times 28$, $\beta=0.44$
Figure 20: Performance comparison between partitioned_contract, SweepContractor [14], CATN [61], and hyperoptimized approximate contraction (HAC) [26] in contracting lattices based on the Ising model. The swap batch size is fixed to be 32 for all experiments. In the legends, “PC” denotes partitioned_contract, MPS/Comb denotes the embedding tree ansatz, and the values (1, 3, 5) denote the size of each partition. In (a)-(f), the number shown on top of each point is the maximum bond size $\chi$. In CATN, “Dmax” is an additional input parameter of the algorithm that controls the size of the MPS uncontracted modes. In HAC, $r$ denotes the distance of the spanning trees used in canonicalization.
Figure 21: Performance comparison between partitioned_contract and SweepContractor [14] in contracting $5\times 5\times 5$ random tensor networks defined on 3D lattices with $\alpha=-0.4$. The swap batch size is fixed to be 32 for all experiments. In the legends, “PC” denotes partitioned_contract, MPS/Comb denotes the embedding tree ansatz, and the values (1, 3, 5) denote the size of each partition. In (a)-(c), the number shown on top of each point is the maximum bond size $\chi$.
@@
(g) Ising, $\beta=0.3$
(h) Ising, $\beta=0.3$
(i) Random, $\alpha=-0.4$
(j) Random, $\alpha=-0.4$
Figure 22: Performance comparison between partitioned_contract and SweepContractor [14] in contracting random regular graphs with degree 3 and 220 vertices. The swap batch size is fixed to be 32 for all experiments. In the legends, “PC” denotes partitioned_contract, MPS/Comb denotes the embedding tree ansatz, and the number (1, 3, 6) denotes the size of each partition. In (a)-(f), the number shown on top of each point is the maximum contracted bond size $\chi$.
@@
We explore the impact of the environment size on the accuracy and efficiency of contracting tensor networks defined on 2D and 3D lattices as well as random regular graphs, and the results are shown in Fig. 20, Fig. 21, and Fig. 22.
In both lattices and random regular graphs, we employ the maximally-unbalanced partial contraction path for the contraction process. This path initiates from one partition and progressively combines the previously-contracted section with a new partition following a linear sequence of the partitions. For 3D lattices, each partition represents either a portion of a fiber or an entire fiber of the lattice. The contraction path is determined through a row- or column-wise traversal of the 2D array resulting from partitioning the 3D lattice into fibers. Regarding random regular graphs, we draw inspiration from [34] to construct the contraction path using a linear ordering of vertices. We achieve this by first employing recursive bisection to generate the linear ordering of all the vertices. Then, we sequentially include a partition consisting of a specified number of tensors into the contraction path, following the order of traversal in the vertex ordering.
When contracting tensor networks defiend on lattices, the results presented in Figs. 20 and 21 reveal that employing a larger parition size (3 or 5 for the $5\times 5\times 5$ 3D grid and 14 for the $28\times 28$ 2D grid) leads to both faster and more accurate contractions when compared to the base condition where each partition contains only one tensor. The improved efficiency arises from using larger partitions, which reduces the number of executions of the density matrix algorithm, offsetting any overhead from using larger environments. Regarding accuracy, we can see from Figs. 20f, 20g, 20j, 20k, 20n and 20o that under the same maximum contracted bond size, utilizing a larger partition size generally yields lower relative errors for both MPS and comb structures. This observation validates the efficacy of the environment in enhancing accuracy. See the next section for a comparison of the MPS and comb structures.
Regarding random regular graphs, the results displayed in Figs. 22b, 22c, 22e, 22f, 22g, 22h, 22i and 22j indicate that using a partition size of 6 results in the best combination of efficiency and accuracy. To summarize, employing a larger partition leads to a larger environment size, generally reducing the contraction error under the same rank. However, when it comes to efficiency, the optimal partition size depends on the specific problem. Factors such as the number of executions of the density matrix algorithm to be performed and the cost of forming the density matrix under different environment sizes need to be taken into consideration to determine the most suitable partition size.
@@
In this section we discuss the relative merits of using the MPS and comb tree ansatzes. When contracting lattices, the results in Figs. 20a, 20b, 20c and 21a demonstrate that both MPS and comb structures exhibit similar performance when the maximum bond size is small. However, as the maximum bond size increases, using the comb ansatz becomes slower in comparison to MPS. On the other hand, when contracting random regular graphs, the results in Figs. 22a and 22d reveal that both structures display similar levels of accuracy and efficiency. In summary, both MPS and comb binary tree structures perform similarly in terms of accuracy. However, the efficiency of the comb structure may lag behind MPS, particularly when dealing with a large rank. This disparity in performance is attributed to the presence of large tensors with size $\chi^{3}$ in the comb ansatz.
@@
Here we compare our proposed partitioned_contract algorithm, along with the CATN algorithm [61], SweepContractor [14], and hyper-optimizated approximate contraction (HAC) [26], on contracting tensor networks defined on lattices and random regular graphs.
As demonstrated in Figs. 20a, 20b, 20c, 21a and 22a, our algorithm consistently outperforms CATN in terms of efficiency across all levels of relative error. When compared to SweepContractor, our algorithm shows superior efficiency for 3D lattices and random graphs, while both methods perform similarly on 2D grids. This is expected, as SweepContractor is specifically optimized for planar graphs. Notably, when contracting a 3D lattice tensor network based on the Ising model, partitioned_contract achieves a 9.2X speed-up compared to both CATN and SweepContractor when reaching a relative error of less than $10^{-9}$. Similarly, when contracting a tensor network with a random regular graph structure based on the Ising model, partitioned_contract achieves a 52.4X speed-up compared to SweepContractor when achieving a relative error of less than $10^{-5}$.
As shown from the results in Figs. 20a, 20b and 20c, HAC is a more efficient baseline compared to CATN and SweepContractor in contracting lattices based on the Ising model. On both $5\times 5\times 5$ and $28\times 28$ grids, both HAC and partitioned_contract can reach pretty high accuracy with similar efficiency. On the $6\times 6\times 6$ grid, the results indicate that partitioned_contract can reach higher accuracy and is more efficient when the relative error is smaller than $10^{-10}$. This can be attributed to HAC generating high-order intermediate tensors, which slows it down and exceeds memory capacity when tested with a rank of 256. In contrast, partitioned_contract performs better due to its use of the MPS ansatz, making it more memory-efficient and effective at higher ranks.
In terms of memory usage, let $s$ denote the size of dimensions in the input graph, $\chi\geq s$ the maximum bond dimension, and $D_{\text{max}}\geq s$ an additional parameter controlling the size of uncontracted MPS modes in CATN. The largest intermediate tensor generated in partitioned_contract has a size of $s\chi^{2}$ for the MPS ansatz and $\chi^{3}$ for the comb ansatz. For CATN, the largest tensor size is $D_{\text{max}}r^{2}$. SweedContractor and partitioned_contract with the MPS ansatz share a similar memory footprint of $s\chi^{2}$. For HAC, the largest tensor size can be $\chi^{d}$ where $d$ is the maximum order of the intermediate tensors. To summarize, partitioned_contract with the MPS ansatz has memory usage that is better than other algorithms as a function of bond dimension.